A manufacturer of light bulbs produces bulbs that last a mean of 950 hours with a standard deviation of 120hours .what is the probability that the mean lifetime of a random sample of 10 of these bulbs is less than 900 hours

1
Expert's answer
2022-03-10T15:40:53-0500

"P(\bar X<900)=P(Z<\frac{900-950}{\frac{120}{\sqrt{10}}})=P(Z<-1.32)=0.0934."